A 65.0-kg worker at a nuclear plant is exposed for 5.00 minutes to 1.00 mCi of beta radiation from 90Sr, each with energy of 0.546 MeV. (a) How many beta radiations does he receive? (b) If 75.0% of the radiation energy is absorbed, what is the total radiation energy (in Joule) absorbed by the worker? What is the radiation dose in gray (Gy) and rad (radiation absorbed dose)? (c) Assuming that the beta radiation has RBE of 0.95, express the radiation dose in milliSievert (mSv) and in millirems (mrem)
Answer #1

1 Ci(Curie) = 3.7 x 1010 radioactive decays per second.

1 mCi(milicurie) = 3.7 x 107 radioactive decays per second.

Energy associated with 1 beta radiation = 0.546 MeV = 8.748 x 10-14J     ( 1 eV = 1.602 x 10-19J)

a) Number of beta radiations he receives in 5 minutes = 5 x 60 x 3.7 x 107 = 1.11 x 1010 beta radiations.

b) Total radiation energy absorbed by the worker = 0.75 x 8.748 x 10-14 x 1.11 x 1010 J = 7.283 x 10-4J.

1 Gray is the absorption of 1 joule of radiation energy per kilogram of matter.

The worker(65 kg) absorbs 7.283 x 10-4 J of energy

The radiation dose is 7.283 x 10-4 J / 65 kg = 1.12 x 10-5 Gy

1 gray = 100 rad.

1.12 x 10-5 Gy = 1.12 x 10-3 rad.

c) RBE gives the effectiveness of the radiation relative to gamma/x ray radiation of the same quantity.

   Effective dose of 1.12 x 10-5 gray beta radiation = 0.95 x 1.12 x 10-5 Sievert. = 1.064 x 10-5 Sv = 1.0.64 x 10-2mSv .

    1 Sv = 100 rem

   1.064 x 10-5 Sv = 1.064 x 10-3 rem = 1.064 mrem